OpenStudy (pagen13):

What is the area of a regular octagon with a side of 8 in.? Round the answer to the nearest tenth? a. 443.4 in. b. 618.0 in. c. 221.7 in. d. 309.0 in.

OpenStudy (anonymous):



OpenStudy (pagen13):

@Jacob can you help?

OpenStudy (anonymous):

if you complete it would be a square so the are would be a square area - 4* triangles area

OpenStudy (anonymous):

* so the area would be square area - 4 triangles area

OpenStudy (anonymous):



OpenStudy (anonymous):

8=x sqrt(2) x=4sqrt(2)

OpenStudy (anonymous):

triangle area= 1/2*b*h=1/2*(4*sqrt(2))*(4sqrt(2))=16

OpenStudy (anonymous):

square area= L*L= (4*sqrt(2)+8+4sqrt(2))^2

OpenStudy (anonymous):

=(8*sqrt(2)+8)^2-4*16=309

OpenStudy (anonymous):

D

OpenStudy (pagen13):

Thank you!